the length of a rectangle is 5 meters longer than the width. find the length and width if the perimeter is 54 meters.
Answer by london maths tutor(243)

Let's length be L and width be w
L = 5 + w
Perimeter = 2(L+w) = 54
2(5+w+w) = 54
2( 5 + 2w) = 54
5 + 2w = 54/2
2w = 27 - 5
w = 22/2
w = 11 m
Therefore, L = 5 + 11 = 16 m
Answer: : Length is 16 m and Width is 11m